# Ergodic behaviors in reversible 3-state cellular automata

### v1

@@ 8 Class IV
Figure 29: Histogram of all dynamical exponents $1/z_{C}$ and $1/z_{D}$ for class IV. The vertical dashed lines, from left to right, represent the values of $1/z\in\{\frac{1}{3},\frac{1}{2},\frac{3}{5},\frac{2}{3}\}$.
@@ 8.2 Subclass IVb
Finally, we note that these models exhibit polynomial decay of at least one correlator with a dynamical exponent $21/z_{C,D}=1/$ 2.777The exception are some $L^{2}$ rules where the correlator can decay exponentially to a small constant value. An example is the automaton with rule 24651387.
@@ 9 Conclusions and outlook
We find a variety of transport phenomena, mostly, but not only with the dynamical exponent $z=2$. They mainly appear in Classes III and IV, and with some also in Class II. Some of these seem to suggest normal diffusive behavior, while others show exceptional (exotic) dynamical structure factors. On top of this we find a plethora of rules with anomalous transport $21>1/z\neq 1/$ 2. We identify possible examples of superdiffusive transport $31/z=2/$ 3 with an unknown (non-KPZ) scaling function. Furthermore, we find a clean example of subdiffusive transport with $31/z=1/$ 3.
@@ Acknowledgements
We would like to thank Matthew Dodelson, Urban Duh, Enej Ilievski, Yusuf Kasim and Luka Paljk for useful discussions. We acknowledge funding from European Research Council (ERC) through Advanced grant QUEST (Grant Agreement No. 101096208, T.P. and R.S.), as well as the Slovenian Research and Innovation agency (ARIS) through the Program P1-0402 (All authors) and Grants N1-0219 (T.P.), N1-0368 (T.P.), N1-0245 (S.G.), J7-60121 (S.G.), and by the STFC Ernest Rutherford Fellowship ST/T00388X/1 (S.G.).
@@ A.1 Basis in the space of extensive local observables
$[\alpha_{1},\alpha_{2}\l\alpha_{r}]_{x}=[\alpha_{1}]_{x}[\alpha_{2}]_{j+1}...[\alpha_{r}]_{x+r-1},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ [\alpha]_{x}(s)=\delta_{\alpha,s_{x}},$ (53)
where $\alpha_{i}\in\{+,-,\varnothing\}$. The observable $[\alpha]_{x}:X^{L}\rightarrow\mathbb{C}$ has the physical meaning of a number of $\alpha$ particles at site $x$.
The local observables add up to a unit element of the algebra $[+]+[-]+[\varnothing]=\mathds{1}$. We want to change the basis, such that the identity is one of the basis elements. We change the basis to an orthogonal one in the following way
$[0]=[\emptyset]+[+]+[-],\quad[1]=[+]-[-],\quad[2]=-2[\emptyset]+[+]+[-],$ (54)
where $[0]\equiv\mathds{1}$. The transformation can be expressed as
@@
Thus, in order to find densities $[\alpha_{1},...\alpha_{r}]$ of observables, that form a basis in $\mathcal{A}_{\mathbb{T}}^{(r)}$, we exclude ones with two or more zeros on the left (except for the identity observable). The dimension of the vector space is then $\dim\left(\mathcal{A}_{\mathbb{T}}^{(r)}\right)=3^{r}-3^{r-2}+1$ and the basis elements are
$A=\sum_{j}\mathbb{T}^{2j}a,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ a=[a_{1}a_{2}...a_{r}]:\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ a_{i}\in\{0,1,2\},\quad a_{1}+a_{2}\neq 0$ (56)
@@
The vector space $\mathcal{A}_{\mathbb{T}}^{(r)}\setminus\mathcal{A}_{\mathbb{T}}^{(r-1)}$ has basis elements with the densities $a=[a_{1}a_{2}...a_{r}],\text{ where}\leavevmode\nobreak\ \leavevmode\nobreak\ (a_{1}+a_{2})\cdot a_{r}\neq 0$; the condition $a_{r}\neq 0$ was imposed for observable not to be in the subspace $\mathcal{A}_{\mathbb{T}}^{(r-1)}$.
@@ A.2 Numerical algorithm
[…]
@@
    Finally, we truncate $b^{\prime}$ to the subspace of observables of range $r$. Namely, we apply $P_{\mathcal{A}^{(r)}_{\mathbb{T}}}[b_{1}\leavevmode\nobreak\ ...\leavevmode\nobreak\ b_{r}b_{r+1}b_{r+2}b_{r+3}]=[b_{1}\leavevmode\nobreak\ ...\leavevmode\nobreak\ b_{r}]$ if $b_{r+1}=b_{r+2}=b_{r+3}=0$ and 00 otherwise. This determines the component of the transfer matrix $\left(\mathcal{T}^{(r)}\right)^{b_{1},...,b_{r}}_{a_{1},...a_{r}}=t^{\prime}_{b_{1},...,b_{r},0,0,0}$.
Repeating these steps over all basis densities $[a_{1}...a_{r}]$ of $\mathcal{A}_{\mathbb{T}}^{(r)}$, we find the matrix representation of the operator $\mathcal{T}^{(r)}$.
@@ Appendix B Detailed properties of 3-state RCA
In this appendix, we present a table of compiled selected data, which is useful for finding interesting models. The rules are ordered according to different classes they belong to and symmetries the exhibit. For each rule, we state the following:
@@
Rule Group Class Return $\kappa$ or $p$ C $\alpha$ or $z_{c}^{-1}$ D $\beta$ or $z_{D}^{-1}$ $|Q|$
@@
38621574 $\mathcal{CPT}$ IIa exp 0.55(2) power 0.538 power 1.126 0 1 1 1 1
38627514 $\mathcal{CPT}$ IIa exp 0.55(2) power 0.537 power 1.131 0 1 1 1 1 1
14725863 $\mathcal{CPT}$ IIa exp 0.56(3) power 0.615 power 0.695 1 1 1 1 1
34267518 $\mathcal{CPT}$ IIa exp 0.55(2) power 0.537 power 1.11 0 1 1 1 1
14625873 $\mathcal{CPT}$ IIa exp 0.55(1) power 0.616 power 0.644 1 1 1 1 1
45721863 $\mathcal{CPT}$ IIa exp 0.56(3) power 0.615 power 0.698 1 1 1 1 1
34268571 $\mathcal{CPT}$ IIa exp 0.54(1) exp 0.014 power 0.65 0 1 1 1 1
58637214 $\mathcal{CPT}$ IIa exp 0.546 exp 0.014 power 0.641 0 1 1 1 1
36478521 $\mathcal{CPT}$ IIa exp 0.7(2) exp 0.096 power 0.535 0 0 1 1 1
36487512 $\mathcal{CPT}$ IIa exp 0.7(2) exp 0.094 power 0.535 0 0 1 1 1
67583142 $\mathcal{CPT}$ IIa exp 0.7(2) exp 0.096 power 0.535 0 0 1 1 1
37186542 $\mathcal{CPT}$ IIa exp 0.6(1) exp 0.094 power 0.535 0 0 1 1 1 1
45621738 $\mathcal{CPT}$ IIa exp 0.9(8) power 0.347 power 0.575 1 2 2 2 2
14625738 $\mathcal{CPT}$ IIa exp 0.9(8) power 0.346 power 0.565 1 2 2 2 2
45721368 $\mathcal{CPT}$ IIa exp 0.55(1) power 0.616 power 0.66 1 1 1 1 1
13465287 $\mathcal{CP}$ IIa exp 0.56(2) power 0.515 power 0.53 0 0 0 0 0 0
35416287 $\mathcal{CP}$ IIa exp 0.9(8) power 0.548 power 0.525 1 1 1 1 1
16435278 $\mathcal{CP+T}$ IIa exp 0.87(8) power 0.46 power 0.413 0 0 0 0 0
56431287 $\mathcal{CP+T}$ IIa exp 0.55(1) exp 0.038 power 0.482 0 0 0 0 0
14625387 $\mathcal{CP+T}$ IIa exp 0.89(9) power 0.411 power 0.534 1 1 1 1 1
14625378 $\mathcal{CP+T}$ IIa exp 0.78(6) power 0.334 power 0.509 2 2 2 2 2
54378621 $\mathcal{CT+P}$ IIa exp 0.56(1) exp 0.038 power 0.522 0 0 0 0 0
45678321 $\mathcal{CT+P}$ IIa exp 0.87(8) power 0.459 power 0.409 0 0 0 0 0
45321678 $\mathcal{CT+P}$ IIa exp 0.89(9) power 0.407 power 0.455 1 1 1 1 1
45621378 $\mathcal{CT+P}$ IIa exp 0.77(5) power 0.333 power 0.505 2 2 2 2 2
46532178 $\mathcal{CP+PT}$ IIa exp 0.79(8) power 0.377 power 0.419 1 1 1 1 1
35216478 $\mathcal{CP+PT}$ IIa exp 0.57(3) power 0.692 power 0.68(3) 2 3 3 3 3
25614378 $\mathcal{CP+PT}$ IIa exp 0.35(3) power 0.32 power 0.386 2 2 2 2 2
12654378 $\mathcal{P+T}$ IIa exp 0.34(3) power 0.319 power 0.392 2 2 2 2 2
12687354 $\mathcal{P+T}$ IIa exp 0.62(4) power 0.692 power 0.69(3) 2 3 3 3 3
34612587 $\mathcal{CPT}$ IIb exp 0.42(6) power 0.501 power 0.501 0 2 2 2 4
14385672 $\mathcal{CPT}$ IIb exp 0.42(5) power 0.501 power 0.5 0 2 2 2 4
34126587 $\mathcal{CP+T}$ IIb exp 0.37(4) power 0.5 power 0.5(3) 0 2 2 2 4
54387612 $\mathcal{CT+P}$ IIb exp 0.39(3) power 0.501 power 0.49(3) 0 2 2 2 4
38167452 $\mathcal{C+T}$ IIb exp 0.39(4) power 0.488 power 0.521 0 2 2 2 4
67438125 $\mathcal{C+T}$ IIb exp 0.41(4) power 0.488 power 0.521 0 2 2 2 4
26534178 $\mathcal{CP+PT}$ IIb exp 0.4(3) power 0.492 power 0.48(3) 2 4 4 4 6
25314678 $\mathcal{CP+PT}$ IIb exp 0.69(3) power 0.567 power 0.628 2 2 4 4 6 6
12654387 $\mathcal{P+T}$ IIb exp 0.69(3) power 0.567 power 0.638 2 2 4 4 6
21678345 $\mathcal{P+T}$ IIb exp 0.39(2) power 0.492 power 0.49(3) 2 4 4 4 6
37268541 $\mathcal{PT+C}$ IIb exp 0.37(2) power 0.487 power 0.52 0 2 2 2 6
68537214 $\mathcal{PT+C}$ IIb exp 0.38(2) power 0.487 power 0.519 0 2 2 2 6
16435287 $\mathcal{CP+T}$ IIc exp 0.36(3) power 0.665 power 0.663 0 2 2 4 4 8
45378621 $\mathcal{CT+P}$ IIc exp 0.36(3) power 0.667 power 0.666 0 2 2 4 4 8
@@
34671528 $\mathcal{CPT}$ IIIb exp 0.78(5) power 0.493 const 0.111 1 2 6 10 18 34
23658471 $\mathcal{CPT}$ IIIb exp 0.26(3) power 0.718 const 0.111 1 2 5 9 17
26457318 $\mathcal{CPT}$ IIIb exp 0.26(3) power 0.718 const 0.111 1 2 5 9 17
56427318 $\mathcal{CPT}$ IIIb exp 0.77(5) power 0.494 const 0.111 1 2 6 10 18
34871526 $\mathcal{CPT}$ IIIb exp 0.78(6) power 0.528 const 0.111 1 2 5 9 17
23758461 $\mathcal{CPT}$ IIIb exp 0.26(3) power 0.506 const 0.111 1 2 6 10 18
56427813 $\mathcal{CPT}$ IIIb exp 0.77(6) power 0.528 const 0.111 1 2 5 9 17
34621578 $\mathcal{CPT}$ IIIb exp 0.8(1) power 0.525 const 0.111 2 3 5 9 17
24657813 $\mathcal{CPT}$ IIIb exp 0.58(3) power 0.517 const 0.111 1 2 4 8 16
54627318 $\mathcal{CPT}$ IIIb exp 0.8(9) power 0.524 const 0.111 2 3 5 9 17
34821576 $\mathcal{CPT}$ IIIb exp 0.9(1) power 0.52 const 0.111 1 2 4 8 16
23751468 $\mathcal{CPT}$ IIIb exp 0.58(4) power 0.517 const 0.111 1 2 4 8 16
54627813 $\mathcal{CPT}$ IIIb exp 0.9(1) power 0.519 const 0.111 1 2 4 8 16
26457813 $\mathcal{CPT}$ IIIb exp 0.26(3) power 0.507 const 0.111 1 2 6 10 18

### v2

@@ 8 Class IV
Figure 31: Histogram of all dynamical exponents $1/z_{C}$ and $1/z_{D}$ for class IV. The vertical dashed lines, from left to right, represent the values of $1/z\in\{\frac{1}{4},\frac{1}{3},\frac{1}{2},\frac{3}{5},\frac{2}{3}\}$.
@@ 8.2 Subclass IVb
Finally, we note that these models exhibit polynomial decay of at least one correlator with a dynamical exponent $1/z_{C,D}=1/2$. [^6] The exception are some $L^{2}$ rules where the correlator can decay exponentially to a small constant value. An example is the automaton with rule 24651387.
@@ 9 Conclusions and outlook
We find a variety of transport phenomena, mostly, but not only with the dynamical exponent $z=2$. They mainly appear in Classes III and IV, and with some also in Class II. Some of these seem to suggest normal diffusive behavior, while others show exceptional (exotic) dynamical structure factors. On top of this we find a plethora of rules with anomalous transport $1>1/z\neq 1/2$. We identify possible examples of superdiffusive transport $1/z=2/3$ with an unknown (non-KPZ) scaling function. Furthermore, we find a clean example of subdiffusive transport with $1/z=1/3$.
@@ Acknowledgements
We would like to thank Luca Delacrétaz, Matthew Dodelson, Urban Duh, Enej Ilievski, Yusuf Kasim and Luka Paljk for useful discussions. We acknowledge funding from European Research Council (ERC) through Advanced grant QUEST (Grant Agreement No. 101096208, T.P. and R.S.), as well as the Slovenian Research and Innovation agency (ARIS) through the Program P1-0402 (All authors) and Grants N1-0219 (T.P.), N1-0368 (T.P.), N1-0245 (S.G.), J7-60121 (S.G.), and by the STFC Ernest Rutherford Fellowship ST/T00388X/1 (S.G.).
@@ A.1 Basis in the space of extensive local observables
$\alpha_{1},\alpha_{2}\l \alpha_{r}]_{x}=[\alpha_{1}]_{x}[\alpha_{2}]_{j+1}...[\alpha_{r}]_{x+r-1},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ [\alpha]_{x}(s)=\delta_{\alpha,s_{x}},$ (53)
where $\alpha_{i}\in\{+,-,\varnothing\}$. The observable $\alpha]_{x}:X^{L}\rightarrow\mathbb{C}$ has the physical meaning of a number of $\alpha$ particles at site $x$.
The local observables add up to a unit element of the algebra $\varnothing]=\mathds{1}$. We want to change the basis, such that the identity is one of the basis elements. We change the basis to an orthogonal one in the following way
$0]=[\emptyset]+[+]+[-],\quad[1]=[+]-[-],\quad[2]=-2[\emptyset]+[+]+[-],$ (54)
where $0]\equiv\mathds{1}$. The transformation can be expressed as
@@
Thus, in order to find densities $\alpha_{1},...\alpha_{r}]$ of observables, that form a basis in $\mathcal{A}_{\mathbb{T}}^{(r)}$, we exclude ones with two or more zeros on the left (except for the identity observable). The dimension of the vector space is then $\dim\left(\mathcal{A}_{\mathbb{T}}^{(r)}\right)=3^{r}-3^{r-2}+1$ and the basis elements are
$A=\sum_{j}\mathbb{T}^{2j}a,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ a=[a_{1}a_{2}...a_{r}]:\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ a_{i}\in\{0,1,2\},\quad a_{1}+a_{2}\neq 0$ (56)
@@
The vector space $\mathcal{A}_{\mathbb{T}}^{(r)}\setminus\mathcal{A}_{\mathbb{T}}^{(r-1)}$ has basis elements with the densities $a=[a_{1}a_{2}...a_{r}],\text{ where}\penalty 10000\ \penalty 10000\ (a_{1}+a_{2})\cdot a_{r}\neq 0$; the condition $a_{r}\neq 0$ was imposed for observable not to be in the subspace $\mathcal{A}_{\mathbb{T}}^{(r-1)}$.
@@ A.2 Numerical algorithm
Since we are interested in the evolution of extensive observables, we change the basis from $\{+,-,\varnothing\}$ to $\{0,1,2\}$. Then, local propagator in new basis can be expressed as
@@
    Finally, we truncate $b^{\prime}$ to the subspace of observables of range $r$. Namely, we apply $P_{\mathcal{A}^{(r)}_{\mathbb{T}}}[b_{1}\penalty 10000\ ...\penalty 10000\ b_{r}b_{r+1}b_{r+2}b_{r+3}]=[b_{1}\penalty 10000\ ...\penalty 10000\ b_{r}]$ if $b_{r+1}=b_{r+2}=b_{r+3}=0$ and 0 otherwise. This determines the component of the transfer matrix $\left(\mathcal{T}^{(r)}\right)^{b_{1},...,b_{r}}_{a_{1},...a_{r}}=t^{\prime}_{b_{1},...,b_{r},0,0,0}$.
Repeating these steps over all basis densities $a_{1}...a_{r}]$ of $\mathcal{A}_{\mathbb{T}}^{(r)}$, we find the matrix representation of the operator $\mathcal{T}^{(r)}$.
@@ Appendix B Detailed properties of 3-state RCA
In this appendix, we present a table of compiled selected data, which is useful for finding interesting models. We note that all the computed data is available at the Zenodo repository. The rules are ordered according to different classes they belong to and symmetries the exhibit. For each rule, we state the following:
@@
Finally, the extended long-time simulations done with systems of size $2^{20}$ are used for the subdiffusive rules 46532178, 16435278, 14625387, 45678321, 45321678, 25614378 and 12654378. The results are presented in Appendix C.
Rule Group Class Return $\kappa/p$ C $1/z_{C}$ D $1/z_{D}$ $|Q|$
@@
38621574 $\mathcal{CPT}$ IIa exp 0.55(2) poly 0.538 poly 1.126 0 1 1 1 1
38627514 $\mathcal{CPT}$ IIa exp 0.55(2) poly 0.537 poly 1.131 0 1 1 1 1 1
14725863 $\mathcal{CPT}$ IIa exp 0.56(3) poly 0.615 poly 0.695 1 1 1 1 1
34267518 $\mathcal{CPT}$ IIa exp 0.55(2) poly 0.537 poly 1.11 0 1 1 1 1
14625873 $\mathcal{CPT}$ IIa exp 0.55(1) poly 0.616 poly 0.644 1 1 1 1 1
45721863 $\mathcal{CPT}$ IIa exp 0.56(3) poly 0.615 poly 0.698 1 1 1 1 1
34268571 $\mathcal{CPT}$ IIa exp 0.54(1) exp 0.014 poly 0.65 0 1 1 1 1
58637214 $\mathcal{CPT}$ IIa exp 0.546 exp 0.014 poly 0.641 0 1 1 1 1
36478521 $\mathcal{CPT}$ IIa exp 0.7(2) exp 0.096 poly 0.535 0 0 1 1 1
36487512 $\mathcal{CPT}$ IIa exp 0.7(2) exp 0.094 poly 0.535 0 0 1 1 1
67583142 $\mathcal{CPT}$ IIa exp 0.7(2) exp 0.096 poly 0.535 0 0 1 1 1
37186542 $\mathcal{CPT}$ IIa exp 0.6(1) exp 0.094 poly 0.535 0 0 1 1 1 1
45621738 $\mathcal{CPT}$ IIa exp 0.9(8) poly 0.347 poly 0.575 1 2 2 2 2
14625738 $\mathcal{CPT}$ IIa exp 0.9(8) poly 0.346 poly 0.565 1 2 2 2 2
45721368 $\mathcal{CPT}$ IIa exp 0.55(1) poly 0.616 poly 0.66 1 1 1 1 1
13465287 $\mathcal{CP}$ IIa exp 0.56(2) poly 0.515 poly 0.53 0 0 0 0 0 0
35416287 $\mathcal{CP}$ IIa exp 0.9(8) poly 0.548 poly 0.525 1 1 1 1 1
16435278 $\mathcal{CP+T}$ IIa exp 0.87(8) poly 0.278 poly 0.265 0 0 0 0 0
56431287 $\mathcal{CP+T}$ IIa exp 0.55(1) exp 0.038 poly 0.482 0 0 0 0 0
14625387 $\mathcal{CP+T}$ IIa exp 0.89(9) poly 0.31(2) poly 0.25(3) 1 1 1 1 1
14625378 $\mathcal{CP+T}$ IIa exp 0.78(6) poly 0.334 poly 0.509 2 2 2 2 2
54378621 $\mathcal{CT+P}$ IIa exp 0.56(1) exp 0.038 poly 0.522 0 0 0 0 0
45678321 $\mathcal{CT+P}$ IIa exp 0.87(8) poly 0.278 poly 0.265 0 0 0 0 0
45321678 $\mathcal{CT+P}$ IIa exp 0.89(9) poly 0.31(2) poly 0.25(3) 1 1 1 1 1
45621378 $\mathcal{CT+P}$ IIa exp 0.77(5) poly 0.333 poly 0.505 2 2 2 2 2
46532178 $\mathcal{CP+PT}$ IIa exp 0.79(8) poly 0.268 poly 0.259 1 1 1 1 1
35216478 $\mathcal{CP+PT}$ IIa exp 0.57(3) poly 0.692 poly 0.68(3) 2 3 3 3 3
25614378 $\mathcal{CP+PT}$ IIa exp 0.35(3) poly 0.27 poly 0.251 2 2 2 2 2
12654378 $\mathcal{P+T}$ IIa exp 0.34(3) poly 0.27 poly 0.251 2 2 2 2 2
12687354 $\mathcal{P+T}$ IIa exp 0.62(4) poly 0.692 poly 0.69(3) 2 3 3 3 3
34612587 $\mathcal{CPT}$ IIb exp 0.42(6) poly 0.501 poly 0.501 0 2 2 2 4
14385672 $\mathcal{CPT}$ IIb exp 0.42(5) poly 0.501 poly 0.5 0 2 2 2 4
34126587 $\mathcal{CP+T}$ IIb exp 0.37(4) poly 0.5 poly 0.5(3) 0 2 2 2 4
54387612 $\mathcal{CT+P}$ IIb exp 0.39(3) poly 0.501 poly 0.49(3) 0 2 2 2 4
38167452 $\mathcal{C+T}$ IIb exp 0.39(4) poly 0.488 poly 0.521 0 2 2 2 4
67438125 $\mathcal{C+T}$ IIb exp 0.41(4) poly 0.488 poly 0.521 0 2 2 2 4
26534178 $\mathcal{CP+PT}$ IIb exp 0.4(3) poly 0.492 poly 0.48(3) 2 4 4 4 6
25314678 $\mathcal{CP+PT}$ IIb exp 0.69(3) poly 0.567 poly 0.628 2 2 4 4 6 6
12654387 $\mathcal{P+T}$ IIb exp 0.69(3) poly 0.567 poly 0.638 2 2 4 4 6
21678345 $\mathcal{P+T}$ IIb exp 0.39(2) poly 0.492 poly 0.49(3) 2 4 4 4 6
37268541 $\mathcal{PT+C}$ IIb exp 0.37(2) poly 0.487 poly 0.52 0 2 2 2 6
68537214 $\mathcal{PT+C}$ IIb exp 0.38(2) poly 0.487 poly 0.519 0 2 2 2 6
16435287 $\mathcal{CP+T}$ IIc exp 0.36(3) poly 0.665 poly 0.663 0 2 2 4 4 8
45378621 $\mathcal{CT+P}$ IIc exp 0.36(3) poly 0.667 poly 0.666 0 2 2 4 4 8
@@
34671528 $\mathcal{CPT}$ IIIb exp 0.78(5) poly 0.493 const 0.111 1 2 6 10 18 34
23658471 $\mathcal{CPT}$ IIIb exp 0.26(3) poly 0.718 const 0.111 1 2 5 9 17
26457318 $\mathcal{CPT}$ IIIb exp 0.26(3) poly 0.718 const 0.111 1 2 5 9 17
56427318 $\mathcal{CPT}$ IIIb exp 0.77(5) poly 0.494 const 0.111 1 2 6 10 18
34871526 $\mathcal{CPT}$ IIIb exp 0.78(6) poly 0.528 const 0.111 1 2 5 9 17
23758461 $\mathcal{CPT}$ IIIb exp 0.26(3) poly 0.506 const 0.111 1 2 6 10 18
56427813 $\mathcal{CPT}$ IIIb exp 0.77(6) poly 0.528 const 0.111 1 2 5 9 17
34621578 $\mathcal{CPT}$ IIIb exp 0.8(1) poly 0.525 const 0.111 2 3 5 9 17
24657813 $\mathcal{CPT}$ IIIb exp 0.58(3) poly 0.517 const 0.111 1 2 4 8 16
54627318 $\mathcal{CPT}$ IIIb exp 0.8(9) poly 0.524 const 0.111 2 3 5 9 17
34821576 $\mathcal{CPT}$ IIIb exp 0.9(1) poly 0.52 const 0.111 1 2 4 8 16
23751468 $\mathcal{CPT}$ IIIb exp 0.58(4) poly 0.517 const 0.111 1 2 4 8 16
54627813 $\mathcal{CPT}$ IIIb exp 0.9(1) poly 0.519 const 0.111 1 2 4 8 16
26457813 $\mathcal{CPT}$ IIIb exp 0.26(3) poly 0.507 const 0.111 1 2 6 10 18